# One-Step Diffusion Samplers via Self-Distillation and Deterministic Flow

### v1

@@ 5.3 Deterministic-Flow Importance Weights
We now evaluate the proposed deterministic-flow (DF) importance weights and compare them against the standard discrete forward-backward path weights (RND). We consider six Bayesian benchmarks from [blessing2024elbos] and the strongly multimodal 40‑mode GMM.
@@
As the number of steps increases, the RND estimator gradually recovers. For NFE $\geq 16$, its $|\Delta\log Z|$ becomes very small, reflecting the asymptotic correctness of the discrete forward-backward likelihood ratio when the integrator is sufficiently fine. However, even in this multi-step regime, RND weights remain high-variance: ESS is consistently lower than for DF, and the resulting ELBOs are also lower. DF, on the other hand, achieves strong ELBOs and ESS while maintaining $|\Delta\log Z|$ below 1 over the entire NFE range.
Overall, these results show that DF importance weights are robust in the one/few-step regime, where discrete forward-backward estimators can catastrophically fail, and remain competitive in the multi-step regime.
5.4 Cost-benefit analysis of one-step inference
@@ Appendix A Notation and Assumptions
We assume that $x\mapsto f(x,t)$ and $x\mapsto u(x,t)$ are globally Lipschitz with at most linear growth, uniformly in $t\in[0,T]$. We assume $\sigma(t)$ is continuous and bounded away from 0 and $\infty$ on $[0,T]$. These standard conditions guarantee strong existence and uniqueness for equations˜2 and 3 [Oksendal2003, Theorem 5.2.1], absolute continuity of the induced path measures, and finiteness of the Radon-Nikodym derivatives we use.
@@
so that the log-Jacobian of $T=\phi_{d}$ along the trajectory starting at $x_{0}$ is $\log|\det\nabla T(x_{0})|=\ell_{d}$ [chen2018ctnf].
@@ B.1 Forward-backward Radon-Nikodym derivative
The result follows the structure of [vargas2024cmcd, proof of Proposition 2.2]. We introduce a reference process with control $w$ and define the corresponding path measures ${{\mathbb{P}}}^{\pi,w}$, $\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}^{\tau,w}$. The Radon-Nikodym derivative can be decomposed as the forward and backward Radon-Nikodym derivatives:
@@
The result then follows by applying Girsanov theorem [nusken2021solving, Section A.1] forward and backward in time. The forward term can be expressed as
@@ B.2 Discrete forward-backward likelihood ratio and a lower bound on $\log Z$
Factorizing the discrete path densities, ${\mathbb{P}}^{\pi,u}({\mathbf{x}}_{0:N})=\pi({\mathbf{x}}_{0})\prod_{n=0}^{N-1}p_{n+1|n}({\mathbf{x}}_{n+1}\mid{\mathbf{x}}_{n})$ and $\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}^{\tau,v}({\mathbf{x}}_{0:N})=\tau({\mathbf{x}}_{N})\prod_{n=0}^{N-1}p_{n|n+1}({\mathbf{x}}_{n}\mid{\mathbf{x}}_{n+1}),$ gives equation˜27 immediately. Taking the expectation of both sides with respect to ${\mathbb{P}}^{\pi,u}$ and noting that $\mathbb{E}_{{\mathbb{P}}^{\pi,u}}\!\bigl[\log\dfrac{{\mathrm{d}}{\mathbb{P}}^{\pi,u}}{{\mathrm{d}}\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}^{\tau,v}}\bigr]=-\mathrm{KL}({\mathbb{P}}^{\pi,u}\|\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}^{\tau,v})\leq 0$ establishes equation˜28. Because the Kullback–Leibler divergence is non‑negative, the bound is always valid, even if $u$ (or $v$) is sub‑optimal or the numerical scheme is coarse. ∎
In discrete time, the Radon-Nikodym derivative between the path measures of a forward simulator and its reverse factorizes into a boundary term and a forward-over-backward product of one-step Markov kernels (equation˜27). Here $p_{n+1|n}$ is the forward one-step kernel of the chosen integrator for the generative SDE, while $p_{n|n+1}$ is the corresponding backward kernel (the conditional of the same discrete chain run backward, i.e., conditioned on the right endpoint). Crucially, $p_{n|n+1}$ is not in general equal to “the EM step of the reverse-time SDE”. See [blessing2025udb, Remark 3.2] for the same discrete RND viewpoint in the underdamped setting.
@@
which is a strict lower bound on $\log Z$ for every $M\geq 1$ and converges monotonically upwards as the control and the time‑step are refined.
@@
Many diffusion papers compute the path log‑weight directly from the Girsanov formula [berner2024dis, richter2024improved],
@@
Discretized as an approximate Riemann sum, this divergence-based Radon-Nikodym derivative does not guarantee a lower-bound for the log-normalization constant [blessing2025udb]. When sampling using $N=1$ steps, the inner running cost of the line-integral discretization vanishes and the estimation collapses to the boundary ratios.
@@ C.2 Hutchinson trace estimator
Evaluating $\nabla\cdot b_{\theta}({\mathbf{x}}_{t},t)=\mathrm{tr}\big(\nabla_{x}b_{\theta}({\mathbf{x}}_{t},t)\big)$ naïvely is $O(d^{2})$ and requires forming Jacobians explicitly. Following [grathwohl2019ffjord], we estimate the trace in linear time using the Hutchinson identity:
@@
Appendix D Implementation Details
D.1 Model architecture and optimization
All methods use a PISGRAD backbone of identical size (same depth/width and Fourier features for time) [zhang2022pis]. Our variant (step-conditioned) augments the backbone with a small embedding for the step size $d$: we compute Fourier features for both $t$ and $d$, concatenate them, and feed the result into the time coder.
We use AdamW [loshchilov2017adamW] with decoupled weight decay $0.1$ and global-norm gradient clipping (threshold $1.0$ by default). All implementations are in JAX/Flax/Optax [jax2018github]. We maintain an EMA of parameters with decay $0.999$. We train and evaluate using a NVIDIA A6000 GPU.
The learning rate is tuned for all methods in $\{0.001,0.0001,0.00001\}$, following [blessing2024elbos]. We use the model checkpoints with the best moving average ELBO.
D.1.1 Jacobian regularization
@@
This regularization mirrors the JVP/trace machinery used in continuous/ODE flows: Hutchinson probes are standard for Jacobian traces in, and controlling Lipschitzness improves numerical stability and reduces stiffness (which otherwise increases solver work). See the instantaneous change-of-variables and trace/JVP discussion in CNFs [chen2018ctnf] and FFJORD [grathwohl2019ffjord].
D.2 Loss algorithmic details
@@
D.3 VP SDE and time schedule
Unless specified otherwise, we use the variance–preserving (VP) noising SDE [ho2020ddpm, song2021sde]

### v2

@@ 5.3 Deterministic-Flow Importance Weights
We now evaluate the proposed deterministic-flow (DF) importance weights and compare them against the standard discrete forward-backward path weights (RND). We consider six Bayesian benchmarks from (Blessing et al., 2024) and the strongly multimodal 40‑mode GMM.
@@
As the number of steps increases, the RND estimator gradually recovers. For NFE $\geq 16$, its $|\Delta\log Z|$ becomes very small, reflecting the asymptotic correctness of the discrete forward-backward likelihood ratio when the integrator is sufficiently fine. However, even in this multi-step regime, RND weights remain high-variance: ESS is consistently lower than for DF, and the resulting ELBOs are also lower. DF, on the other hand, achieves strong ELBOs and ESS while maintaining $|\Delta\log Z|$ below 1 over the entire NFE range. Overall, these results show that DF importance weights are robust in the one/few-step regime, where discrete forward-backward estimators can catastrophically fail, and remain competitive in the multi-step regime.
@@ CONCLUSION
ACKNOWLEDGEMENTS
This research is supported in part by NSF IIS-2508145 and Amazon Research Award.
@@ Appendix A Notation and Assumptions
We assume that $x\mapsto f(x,t)$ and $x\mapsto u(x,t)$ are globally Lipschitz with at most linear growth, uniformly in $t\in[0,T]$. We assume $\sigma(t)$ is continuous and bounded away from 0 and $\infty$ on $[0,T]$. These standard conditions guarantee strong existence and uniqueness for equations˜2 and 3 (Øksendal, 2003, Theorem 5.2.1), absolute continuity of the induced path measures, and finiteness of the Radon-Nikodym derivatives we use.
@@
so that the log-Jacobian of $T=\phi_{d}$ along the trajectory starting at $x_{0}$ is $\log|\det\nabla T(x_{0})|=\ell_{d}$ (Chen et al., 2018a).
@@ B.1 Forward-backward Radon-Nikodym derivative
The result follows the structure of (Vargas et al., 2024, proof of Proposition 2.2). We introduce a reference process with control $w$ and define the corresponding path measures ${{\mathbb{P}}}^{\pi,w}$, $\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}^{\tau,w}$. The Radon-Nikodym derivative can be decomposed as the forward and backward Radon-Nikodym derivatives:
@@
The result then follows by applying Girsanov theorem (Nüsken and Richter, 2021, Section A.1) forward and backward in time. The forward term can be expressed as
@@ B.2 Discrete forward-backward likelihood ratio and a lower bound on $\log Z$
Factorizing the discrete path densities, ${\mathbb{P}}^{\pi,u}({\mathbf{x}}_{0:N})=\pi({\mathbf{x}}_{0})\prod_{n=0}^{N-1}p_{n+1|n}({\mathbf{x}}_{n+1}\mid{\mathbf{x}}_{n})$ and $\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}^{\tau,v}({\mathbf{x}}_{0:N})=\tau({\mathbf{x}}_{N})\prod_{n=0}^{N-1}p_{n|n+1}({\mathbf{x}}_{n}\mid{\mathbf{x}}_{n+1}),$ gives equation˜27 immediately. Taking the expectation of both sides with respect to ${\mathbb{P}}^{\pi,u}$ and noting that $\mathbb{E}_{{\mathbb{P}}^{\pi,u}}\bigl[\log\dfrac{{\mathrm{d}}{\mathbb{P}}^{\pi,u}}{{\mathrm{d}}\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}^{\tau,v}}\bigr]=\mathrm{KL}({\mathbb{P}}^{\pi,u}\|\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}^{\tau,v})\geq 0$ establishes equation˜28. Because the Kullback–Leibler divergence is non‑negative, the bound is always valid, even if $u$ (or $v$) is sub‑optimal or the numerical scheme is coarse. ∎
In discrete time, the Radon-Nikodym derivative between the path measures of a forward simulator and its reverse factorizes into a boundary term and a forward-over-backward product of one-step Markov kernels (equation˜27). Here $p_{n+1|n}$ is the forward one-step kernel of the chosen integrator for the generative SDE, while $p_{n|n+1}$ is the corresponding backward kernel (the conditional of the same discrete chain run backward, i.e., conditioned on the right endpoint). Crucially, $p_{n|n+1}$ is not in general equal to “the EM step of the reverse-time SDE”. See (Blessing et al., 2025, Remark 3.2) for the same discrete RND viewpoint in the underdamped setting.
@@
Drawing $M$ paths yields an unbiased estimator of the ELBO (the LHS of equation 28). Since the ELBO lower-bounds $\log Z$, this provides a lower bound in expectation.
@@
Many diffusion papers compute the path log‑weight directly from the Girsanov formula (Berner et al., 2024; Richter and Berner, 2024),
@@
Discretized as an approximate Riemann sum, this divergence-based Radon-Nikodym derivative does not guarantee a lower-bound for the log-normalization constant (Blessing et al., 2025). When sampling using $N=1$ steps, the inner running cost of the line-integral discretization vanishes and the estimation collapses to the boundary ratios.
@@ C.2 Hutchinson trace estimator
Evaluating $\nabla\cdot b_{\theta}({\mathbf{x}}_{t},t)=\mathrm{tr}\big(\nabla_{x}b_{\theta}({\mathbf{x}}_{t},t)\big)$ naïvely is $O(d^{2})$ and requires forming Jacobians explicitly. Following Grathwohl et al. (2019), we estimate the trace in linear time using the Hutchinson identity:
@@
Appendix D Cost-benefit analysis of one-step inference
Appendix E Implementation Details
E.1 Model architecture and optimization
All methods use a PISGRAD backbone of identical size (same depth/width and Fourier features for time) (Zhang and Chen, 2022). Our variant (step-conditioned) augments the backbone with a small embedding for the step size $d$: we compute Fourier features for both $t$ and $d$, concatenate them, and feed the result into the time coder.
We use AdamW (Loshchilov and Hutter, 2017) with decoupled weight decay $0.1$ and global-norm gradient clipping (threshold $1.0$ by default). All implementations are in JAX/Flax/Optax (Bradbury et al., 2018). We maintain an EMA of parameters with decay $0.999$. We train and evaluate using a NVIDIA A6000 GPU.
The learning rate is tuned for all methods in $\{0.001,0.0001,0.00001\}$, following Blessing et al. (2024). We use the model checkpoints with the best moving average ELBO.
E.1.1 Jacobian regularization
@@
This regularization mirrors the JVP/trace machinery used in continuous/ODE flows: Hutchinson probes are standard for Jacobian traces in, and controlling Lipschitzness improves numerical stability and reduces stiffness (which otherwise increases solver work). See the instantaneous change-of-variables and trace/JVP discussion in CNFs (Chen et al., 2018a) and FFJORD (Grathwohl et al., 2019).
E.2 Loss algorithmic details
@@
E.3 VP SDE and time schedule
Unless specified otherwise, we use the variance–preserving (VP) noising SDE (Ho et al., 2020; Song et al., 2021)